OpenStudy (anonymous):

if a compound is 50.7% C, 4.2% H and 45.1% O and the molar mass is 142 g/mol what is the molecular formula?

OpenStudy (somy):

well firstly when it comes to empirical formula, in the questions its quite usual to give percentages and these percentages are considered to be 'masses' so take all percentages as masses of each element and then find mole of each element \[mole= \frac{ mass }{ Mr }\] Mr is molecular mass and you can find them from periodic table i'll show you how to find mole of one of the given elements, the rest you try doing yourself Im gonna find mole for C \[mole=\frac{ 50.7 }{ 12 }\] = 4.225 moles repeat same way for the other two DO NOT round moles of, leave them with as many decimals as they are afterwards you'll get moles for each element, and now look which mole value is the SMALLEST between those three values then divide all the 3 moles by that SMALLEST mole now you well get 3 new moles, if ANY mole after division is like this 1.67 for example DO NOT ROUND IT OF what you should to is try multiplying it to a number that will not give you decimals try multiplying by 2 = 2* 1.67= 3.34 nah this one is no good now by 3= 3* 1.67= 5.01 now this one is good, because its very close to a full number so you can take it as 5, but of course if you are getting full number with no decimals it good, so that will be your final mole of that element if you multiplied that mole of that element by 3 as i did in example, you have to multiply the other 2 moles also by 3 and then you will get moles for each element for example you got C= 1 mole, H= 4 moles, and O- 1 moles your empirical formula will look like this \[C _{1}H _{4}O _{1}\] meaning \[C H _{4}O \] now find the total Mr of this empirical formula which is = 12+4+16= 32 then take the Mr given for the molecular formula and divide it by Mr of your empirical formula, lets suppose you did it and got 2 or 2.23 here you can round of so if its 2.23 then just take it as 2 so that'll mean \[[CH _{4}O] \times2 = C _{2}H _{8}O _{2}\] and this will be the Molecular formula so do all the working of your question the way i did in examples I gave you, and you will be able to find molecular formula of the given question :) if you have any questions feel free to ask :D Hope that helps!

OpenStudy (somy):

please do not confuse my working with your question because i did not solve your question, i gave you examples of HOW to solve it so that you can do it on your own :D